### v1

@@ 1.3 Previous results
To prove (conditional) exponential lower bounds, the standard assumptions that $\mathsf{P}\not=\mathsf{NP}$ or $\mathsf{VP}\not=\mathsf{VNP}$ are not enough. It is consistent with our current knowledge that for instance $\mathsf{P}\not=\mathsf{NP}$, but $\mathsf{NP}$-hard problems can have subexponential time algorithms. What we need is a complexity assumption stating that certain problems can only be solved in exponential time. This is the exponential time hypothesis (ETH) in the Boolean setting. Dell et al. [11] studied the exponential time complexity of the permanent, they prove that when there is an algorithm for computing the permanent in time $2^{o(n)}$, then this violates the counting version of the exponential time hypothesis #ETH. #ETH states that there is a constant $c$ such that no deterministic algorithm can count the number of satisfying assignments of a formula in $3$-CNF in time $2^{cn}$. For connections between parameterized and subexponential complexity in the Boolean setting, we refer to [13, 14].
Bläser and Engels [4] transfer the important definitions and results from parameterized complexity in the Boolean world to define a theory of parameterized algebraic complexity classes. In particular, they define the $\mathsf{VW}$-hierarchy and prove that the clique polynomial and the $k$-permanent are $\mathsf{VW}[1]$-complete (under so-called fpt-substitutions). They also claim the hardness of the restricted permanent for the class $\mathsf{VW}[t]$ for every constant $t$ and sketch a proof. Note that $\mathsf{VW}[\mathsf{F}]$ contains each $\mathsf{VW}[t]$. So we strengthen the hardness proof in [4] and complement it with an upper bound.
The main tool used by Bürgisser [9] to prove the results above is the counting hierarchy. The polynomial counting hierarchy was introduced by Wagner [29] to classify the complexity of Boolean counting problems. The fact that small circuits for the permanent collapses the counting hierarchy is used by Bürgisser to prove the results mentioned above.
Finally, there have been quite a few works [9, 18, 20, 19], where we have conditional separations on the constant-free version of $\mathsf{VP}$ and $\mathsf{VNP}$, namely $\mathsf{VP}^{0}$ and $\mathsf{VNP}^{0}$, or their variants, depending on the strength of the conjecture. But this is the first time that we are separating algebraic classes and proving exponential lower bounds, assuming the $\tau$-conjecture.
@@ 1.4 Structure of the paper
In Section˜2, we defined the basics of constant-free Valiant’s model and the unbounded and parameterized setting. In Section˜3, we introduce the linear counting hierarchy ($\mathsf{CH}_{\mathrm{lin}}$) and its basic properties. Section˜4 connects Valiant’s model to the counting hierarchy. Here, we formally introduce exponential sums and investigate their relation to the parameterized classes. The main result is that the fixed-parameter tractability of exponential sums collapses the counting hierarchy. The proofs are quite similar to [9], however, we need to pay special attention to the fact the witness size is linear. Section˜5 introduces the definability (computability) of integers in the linear counting hierarchy, and some closure properties of the same. Section˜6 proves the exponential lower bound on exponential sum assuming $\tau$-conjecture. Section˜7 introduces the parameterized $\mathsf{VW}$-classes and its basic properties. In Section 8 we prove some easy conditional collapse results of the $\mathsf{VW}$-hierarchy in various circuit models.
@@ 2.1 Constant-free and unbounded models
It is not clear whether showing $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$ implies $\mathsf{VP}\neq\mathsf{VNP}$, it is not even clear whether $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$ $\implies$ $\tau(\operatorname{per}_{n})=n^{\omega(1)}$. The subtlety here is that in the algebraic completeness proof for the permanent, divisions by two occur! However, a partial implication is known due to [9, Theorem 2.10]: Showing $\tau(2^{p(n)}f_{n})=n^{\omega(1)}$, for some $f_{n}\in\mathsf{VNP}^{0}$ and all $p\text{-}\mathsf{bounded}p(n)$ would imply that $\tau(\operatorname{per}_{n})=n^{\omega(1)}$.
@@
Valiant’s classes in the unbounded setting: It is well-known that an algebraic circuit of size $s$, can compute polynomials of degree $\exp(s)$; e.g., $f(x)=x^{2^{s}}$, and $L(f)=O(s)$. This brings us to the next definition, the class $\mathsf{VP}_{\textup{nb}}$, originally defined in [21]. A sequence of polynomials $(f)=(f_{n})_{n}\in\mathsf{VP}_{\textup{nb}}$, if the number of variables in $f_{n}$ and $L(f_{n})$ are both $p\text{-}\mathsf{bounded}$ (the degree may be exponentially large). The subscript “nb” signifies the “not bounded” phenomenon on the degree of the polynomial, in contrast to the original class $\mathsf{VP}$. Similarly, a sequence of polynomials $(f)=(f_{n})_{n}\in\mathsf{VNP}_{\textup{nb}}$, if there exists a $p\text{-}\mathsf{bounded}$ function $q(n)$ and $g_{n}(\mathbf{X},Y_{1},\dots,Y_{q(n)})\in\mathsf{VP}_{\textup{nb}}$ where
@@
One can analogously define $\mathsf{VP}_{\textup{nb}}^{0}$ and $\mathsf{VNP}_{\textup{nb}}^{0}$, in the constant-free setting. It is obvious that $\mathsf{VP}_{\textup{nb}}=\mathsf{VNP}_{\textup{nb}}$ implies $\mathsf{VP}=\mathsf{VNP}$, but the converse is unclear. However, [21] showed that over a ring of positive characteristic, the converse holds, i.e., $\mathsf{VP}=\mathsf{VNP}$ implies $\mathsf{VP}_{\textup{nb}}=\mathsf{VNP}_{\textup{nb}}$! On the other hand, [20] showed that $\mathsf{VP}^{0}=\mathsf{VNP}^{0}$ implies that $\mathsf{VP}_{\textup{nb}}^{0}=\mathsf{VNP}_{\textup{nb}}^{0}$, and the converse is unclear because it seems difficult to rule out the possibility that some polynomial family in $\mathsf{VNP}^{0}$ does not lie in $\mathsf{VP}^{0}$, but still in $\mathsf{VP}$ (i.e., computable by polynomial-size algebraic circuits using exponentially large-bit integers).
@@ 2.2 Parameterized Valiant’s classes
Parameterized Valiant’s classes were introduced in [4]. We will briefly review the definitions and results there and extend them to the constant-free and unbounded setting. We first start with the fixed-parameter tractable classes. The $W$-hierarchies will be introduced later since we only need them in the second part of this work.
@@
The most natural parameterization is by the degree: Let $(p_{n})$ be any $p\text{-}\mathsf{family}$ then we get a parameterized family $(p_{n,k})$ by setting $p_{n,k}:=$ the homogeneous part of degree $k$ of $p_{n}$. For more details, we will refer the reader to [4].
@@
However p-projection in Valiant’s world seems to be weaker compared to parsimonious poly-time reduction in the Boolean world; therefore we need a stronger notion of reduction for defining algebraic models of the Boolean #W\#W-classes, see [4]. That’s why we are defining substitutions. We will analogously define it for constant-free model as well.
@@
One can define constant-free fpt-projections analogously. The following lemma should be immediate from the definitions, see [4] for a proof in the case of $\mathsf{VFPT}$.
@@ 3 Linear counting hierarchy
In this section, we define the linear counting hierarchy, a variant of the counting hierarchy, which will allow us to talk about subexponential complexity. The original counting hierarchy was defined by Wagner [29]. We here restrict the witness length to be linear, which is important when dealing with exponential complexity. Allender et al. [2] also define a linear counting hierarchy. Their definition is not comparable to ours. We use an operator-based definition: The base class is deterministic polynomial time and the witness length is linearly bounded. Allender et al. use an oracle TM definition: The oracle Turing machine is probabilistic and linear time bounded, which automatically bounds the query lengths.
@@
It is clear that $\mathbf{C}^{\prime}_{\mathsf{lin}}.K\subseteq\mathbf{C}_{\mathsf{lin}}.K$ for any class $K$. Moreover, by an easy adaption of the proof of [26, Lemma 3.3], for any language $K\in\mathsf{CH}$, $\mathbf{C}_{\mathsf{lin}}.K\subseteq\mathbf{C}^{\prime}_{\mathsf{lin}}.K$. Also, from the definition, we can say that $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}\subseteq\mathsf{CH}$. Therefore, the following holds.
@@ 5 Integers definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$
In [9, Section 3], integers are studied that are definable in the counting hierarchy. We adapt this notation to the linear counting hierarchy. Formally, we are given a sequence of integers $(a(n,k))_{n\in\mathbb{N},k\leq q(n)}$ for some $p\text{-}\mathsf{bounded}$ function $q:\mathbb{N}\rightarrow\mathbb{N}$. We can assume that $|a(n,k)|\leq 2^{n^{c}}$ for some constant $c$. In other words, the bit-size of $a(n,k)$ is at most exponential, as we think $n,k$ has been represented in binary by $O(\log n)$ bits. Now consider two languages,
@@
Now we show an essential criterion for a sequence to be in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. It is an adaption with some additional modifications and observations from [16], which were further implemented in [9, Theorem 3.5].
@@
Our argument goes similar to [9, Theorem 3.5], with some further modifications.
At first, let us show that for nonnegative sequences, $(a)$ is definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}\iff\mathsf{CR}(a)\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. To show the $\Rightarrow$ direction, start with a Dlogtime-uniform $\mathsf{TC}^{0}$ circuit family $(\mathcal{C}_{n})_{n}$ which computes the Chinese remainder representation of an $n$-bit number, modulo primes $p<n^{2}$, from its binary representation. By [16, Lemma 4.1], $\mathcal{C}_{n}$ has size ${\rm poly}(n)$ and constant depth $D$. Consider the language
@@
for $d\in\{0,\dots,D\}$ and $(n,k,F)$ are given by their binary encoding. [9, Theorem 3.5] shows that $L_{d+1}\in\mathbf{C}^{\prime}.L_{d}$. But in fact we can say even stronger that $L_{d+1}\in\mathbf{C}_{\mathsf{lin}}^{\prime}.{L_{d}}$. This is because when we are given $(n,k,F,b)$ as input by their binary encoding and $F$ is some majority gate at depth $d+1$, we need to check if $(n,k,G,1)$ is in $L_{d}$ for all gates $G$ at depth $d$ connected to $F$. The lengths of the witnesses $(n,k,G,1)$ is $O(\log n)$, which is linear in the input. By Dlogtime-uniformity of $(\mathcal{C}_{n})_{n}$, we can check if $G$ is connected to $F$ in polynomial time. And computing the majority of at most ${\rm poly}(n)$ many gates can be done by checking
@@
for some linear function $\ell$. Hence, our claim is true. The rest of the proof and the other direction is similar to the argument given in [9]. Hence, $(a)$ is definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}\iff\mathsf{CR}(a)\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$.
@@
The proof is again similar to [9, Theorem 3.10].
Part (i): $(b(n))_{n\in\mathbb{N}}\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. By [28], we know that iterated addition of $n$ many numbers $0\leq X_{1},\dots,X_{n}\leq 2^{n}$, given in their binary representation, can be done by Dlogtime-uniform $\mathsf{TC}^{0}$ circuits. Say this circuit family is $(\mathcal{C}_{n})_{n}$. $\mathcal{C}_{n}$ has ${\rm poly}(n)$ size and constant depth $D$. Now, we can take some $\mathcal{C}_{n^{c^{\prime}}}$ for some suitable constant $c^{\prime}$ and using the idea same as in Theorem 5.2, we can say that $(b(n):=\sum_{k=0}^{q(n)}a(n,k))_{n\in\mathbb{N}}$ is definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. Note that while we are summing a polynomial number of numbers, the bit-size for addressing the elements of $a(n,k)$ is $\log n+\log k=O(\log n)$, which is linear in the input size.

### v2

@@ 1.3 Previous results
To prove (conditional) exponential lower bounds, the standard assumptions that $\mathsf{P}\not=\mathsf{NP}$ or $\mathsf{VP}\not=\mathsf{VNP}$ are not enough. It is consistent with our current knowledge that for instance $\mathsf{P}\not=\mathsf{NP}$, but $\mathsf{NP}$-hard problems can have subexponential time algorithms. What we need is a complexity assumption stating that certain problems can only be solved in exponential time. This is the exponential time hypothesis (ETH) in the Boolean setting. Dell et al. [DHM+14] studied the exponential time complexity of the permanent, they prove that when there is an algorithm for computing the permanent in time $2^{o(n)}$, then this violates the counting version of the exponential time hypothesis #ETH. #ETH states that there is a constant $c$ such that no deterministic algorithm can count the number of satisfying assignments of a formula in $3$-CNF in time $2^{cn}$. For connections between parameterized and subexponential complexity in the Boolean setting, we refer to [FG04, FG06].
Bläser and Engels [BE19] transfer the important definitions and results from parameterized complexity in the Boolean world to define a theory of parameterized algebraic complexity classes. In particular, they define the $\mathsf{VW}$-hierarchy and prove that the clique polynomial and the $k$-permanent are $\mathsf{VW}[1]$-complete (under so-called fpt-substitutions). They also claim the hardness of the restricted permanent for the class $\mathsf{VW}[t]$ for every constant $t$ and sketch a proof. Note that $\mathsf{VW}[\mathsf{F}]$ contains each $\mathsf{VW}[t]$. So we strengthen the hardness proof in [BE19] and complement it with an upper bound.
The main tool used by Bürgisser [BÜR09] to prove the results above is the counting hierarchy. The polynomial counting hierarchy was introduced by Wagner [WAG86] to classify the complexity of Boolean counting problems. The fact that small circuits for the permanent collapses the counting hierarchy is used by Bürgisser to prove the results mentioned above.
Finally, there have been quite a few works [BÜR09, KOI05, KP11, KOI11], where we have conditional separations on the constant-free version of $\mathsf{VP}$ and $\mathsf{VNP}$, namely $\mathsf{VP}^{0}$ and $\mathsf{VNP}^{0}$, or their variants, depending on the strength of the conjecture. But this is the first time that we are separating algebraic classes and proving exponential lower bounds, assuming the $\tau$-conjecture.
@@ 1.4 Structure of the paper
In Section˜2, we defined the basics of constant-free Valiant’s model and the unbounded and parameterized setting. In Section˜3, we introduce the linear counting hierarchy ($\mathsf{CH}_{\mathrm{lin}}$) and its basic properties. Section˜4 connects Valiant’s model to the counting hierarchy. Here, we formally introduce exponential sums and investigate their relation to the parameterized classes. The main result is that the fixed-parameter tractability of exponential sums collapses the counting hierarchy. The proofs are quite similar to [BÜR09], however, we need to pay special attention to the fact the witness size is linear. Section˜5 introduces the definability (computability) of integers in the linear counting hierarchy, and some closure properties of the same. Section˜6 proves the exponential lower bound on exponential sum assuming $\tau$-conjecture. Section˜7 introduces the parameterized $\mathsf{VW}$-classes and its basic properties. In Section 8 we prove some easy conditional collapse results of the $\mathsf{VW}$-hierarchy in various circuit models.
@@ 2.1 Constant-free and unbounded models
It is not clear whether showing $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$ implies $\mathsf{VP}\neq\mathsf{VNP}$, it is not even clear whether $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$ $\implies$ $\tau(\operatorname{per}_{n})=n^{\omega(1)}$. The subtlety here is that in the algebraic completeness proof for the permanent, divisions by two occur! However, a partial implication is known due to [BÜR09, Theorem 2.10]: Showing $\tau(2^{p(n)}f_{n})=n^{\omega(1)}$, for some $f_{n}\in\mathsf{VNP}^{0}$ and all $p\text{-}\mathsf{bounded}p(n)$ would imply that $\tau(\operatorname{per}_{n})=n^{\omega(1)}$.
@@
Valiant’s classes in the unbounded setting: It is well-known that an algebraic circuit of size $s$, can compute polynomials of degree $\exp(s)$; e.g., $f(x)=x^{2^{s}}$, and $L(f)=O(s)$. This brings us to the next definition, the class $\mathsf{VP}_{\textup{nb}}$, originally defined in [MAL07]. A sequence of polynomials $(f)=(f_{n})_{n}\in\mathsf{VP}_{\textup{nb}}$, if the number of variables in $f_{n}$ and $L(f_{n})$ are both $p\text{-}\mathsf{bounded}$ (the degree may be exponentially large). The subscript “nb” signifies the “not bounded” phenomenon on the degree of the polynomial, in contrast to the original class $\mathsf{VP}$. Similarly, a sequence of polynomials $(f)=(f_{n})_{n}\in\mathsf{VNP}_{\textup{nb}}$, if there exists a $p\text{-}\mathsf{bounded}$ function $q(n)$ and $g_{n}(\mathbf{X},Y_{1},\dots,Y_{q(n)})\in\mathsf{VP}_{\textup{nb}}$ where
@@
One can analogously define $\mathsf{VP}_{\textup{nb}}^{0}$ and $\mathsf{VNP}_{\textup{nb}}^{0}$, in the constant-free setting. It is obvious that $\mathsf{VP}_{\textup{nb}}=\mathsf{VNP}_{\textup{nb}}$ implies $\mathsf{VP}=\mathsf{VNP}$, but the converse is unclear. However, [MAL07] showed that over a ring of positive characteristic, the converse holds, i.e., $\mathsf{VP}=\mathsf{VNP}$ implies $\mathsf{VP}_{\textup{nb}}=\mathsf{VNP}_{\textup{nb}}$! On the other hand, [KP11] showed that $\mathsf{VP}^{0}=\mathsf{VNP}^{0}$ implies that $\mathsf{VP}_{\textup{nb}}^{0}=\mathsf{VNP}_{\textup{nb}}^{0}$, and the converse is unclear because it seems difficult to rule out the possibility that some polynomial family in $\mathsf{VNP}^{0}$ does not lie in $\mathsf{VP}^{0}$, but still in $\mathsf{VP}$ (i.e., computable by polynomial-size algebraic circuits using exponentially large-bit integers).
@@ 2.2 Parameterized Valiant’s classes
Parameterized Valiant’s classes were introduced in [BE19]. We will briefly review the definitions and results there and extend them to the constant-free and unbounded setting. We first start with the fixed-parameter tractable classes. The $W$-hierarchies will be introduced later since we only need them in the second part of this work.
@@
The most natural parameterization is by the degree: Let $(p_{n})$ be any $p\text{-}\mathsf{family}$ then we get a parameterized family $(p_{n,k})$ by setting $p_{n,k}:=$ the homogeneous part of degree $k$ of $p_{n}$. For more details, we will refer the reader to [BE19].
@@
However p-projection in Valiant’s world seems to be weaker compared to parsimonious poly-time reduction in the Boolean world; therefore we need a stronger notion of reduction for defining algebraic models of the Boolean #W\#W-classes, see [BE19]. That’s why we are defining substitutions. We will analogously define it for constant-free model as well.
@@
One can define constant-free fpt-projections analogously. The following lemma should be immediate from the definitions, see [BE19] for a proof in the case of $\mathsf{VFPT}$.
@@ 3 Linear counting hierarchy
In this section, we define the linear counting hierarchy, a variant of the counting hierarchy, which will allow us to talk about subexponential complexity. The original counting hierarchy was defined by Wagner [WAG86]. We here restrict the witness length to be linear, which is important when dealing with exponential complexity. Allender et al. [AKR+01] also define a linear counting hierarchy. Their definition is not comparable to ours. We use an operator-based definition: The base class is deterministic polynomial time and the witness length is linearly bounded. Allender et al. use an oracle TM definition: The oracle Turing machine is probabilistic and linear time bounded, which automatically bounds the query lengths.
@@
It is clear that $\mathbf{C}^{\prime}_{\mathsf{lin}}.K\subseteq\mathbf{C}_{\mathsf{lin}}.K$ for any class $K$. Moreover, by an easy adaption of the proof of [TOR91, Lemma 3.3], for any language $K\in\mathsf{CH}$, $\mathbf{C}_{\mathsf{lin}}.K\subseteq\mathbf{C}^{\prime}_{\mathsf{lin}}.K$. Also, from the definition, we can say that $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}\subseteq\mathsf{CH}$. Therefore, the following holds.
@@ 5 Integers definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$
In [BÜR09, Section 3], integers are studied that are definable in the counting hierarchy. We adapt this notation to the linear counting hierarchy. Formally, we are given a sequence of integers $(a(n,k))_{n\in\mathbb{N},k\leq q(n)}$ for some $p\text{-}\mathsf{bounded}$ function $q:\mathbb{N}\rightarrow\mathbb{N}$. We can assume that $|a(n,k)|\leq 2^{n^{c}}$ for some constant $c$. In other words, the bit-size of $a(n,k)$ is at most exponential, as we think $n,k$ has been represented in binary by $O(\log n)$ bits. Now consider two languages,
@@
Now we show an essential criterion for a sequence to be in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. It is an adaption with some additional modifications and observations from [HAB02], which were further implemented in [BÜR09, Theorem 3.5].
@@
Our argument goes similar to [BÜR09, Theorem 3.5], with some further modifications.
At first, let us show that for nonnegative sequences, $(a)$ is definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}\iff\mathsf{CR}(a)\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. To show the $\Rightarrow$ direction, start with a Dlogtime-uniform $\mathsf{TC}^{0}$ circuit family $(\mathcal{C}_{n})_{n}$ which computes the Chinese remainder representation of an $n$-bit number, modulo primes $p<n^{2}$, from its binary representation. By [HAB02, Lemma 4.1], $\mathcal{C}_{n}$ has size ${\rm poly}(n)$ and constant depth $D$. Consider the language
@@
for $d\in\{0,\dots,D\}$ and $(n,k,F)$ are given by their binary encoding. [BÜR09, Theorem 3.5] shows that $L_{d+1}\in\mathbf{C}^{\prime}.L_{d}$. But in fact we can say even stronger that $L_{d+1}\in\mathbf{C}_{\mathsf{lin}}^{\prime}.{L_{d}}$. This is because when we are given $(n,k,F,b)$ as input by their binary encoding and $F$ is some majority gate at depth $d+1$, we need to check if $(n,k,G,1)$ is in $L_{d}$ for all gates $G$ at depth $d$ connected to $F$. The lengths of the witnesses $(n,k,G,1)$ is $O(\log n)$, which is linear in the input. By Dlogtime-uniformity of $(\mathcal{C}_{n})_{n}$, we can check if $G$ is connected to $F$ in polynomial time. And computing the majority of at most ${\rm poly}(n)$ many gates can be done by checking
@@
for some linear function $\ell$. Hence, our claim is true. The rest of the proof and the other direction is similar to the argument given in [BÜR09]. Hence, $(a)$ is definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}\iff\mathsf{CR}(a)\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$.
@@
The proof is again similar to [BÜR09, Theorem 3.10].
Part (i): $(b(n))_{n\in\mathbb{N}}\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. By [VOL99], we know that iterated addition of $n$ many numbers $0\leq X_{1},\dots,X_{n}\leq 2^{n}$, given in their binary representation, can be done by Dlogtime-uniform $\mathsf{TC}^{0}$ circuits. Say this circuit family is $(\mathcal{C}_{n})_{n}$. $\mathcal{C}_{n}$ has ${\rm poly}(n)$ size and constant depth $D$. Now, we can take some $\mathcal{C}_{n^{c^{\prime}}}$ for some suitable constant $c^{\prime}$ and using the idea same as in Theorem 5.2, we can say that $(b(n):=\sum_{k=0}^{q(n)}a(n,k))_{n\in\mathbb{N}}$ is definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. Note that while we are summing a polynomial number of numbers, the bit-size for addressing the elements of $a(n,k)$ is $\log n+\log k=O(\log n)$, which is linear in the input size.